Ten identical charges of amount 

7

 

μ

C

7 μC

 each are closed in a rectangular box. What is the electric flux through the box?


1
Expert's answer
2020-07-14T08:52:11-0400

According to the Gauss's law, the electric flux "\Phi_E" through the box is proportional to the total charge "Q" enclosed within this box:


"\Phi_E = \dfrac{Q}{\varepsilon_0}"

where "\varepsilon_0 = 8.85\times10^{-12}F/m".

The total charge is


"Q = 10\cdot 7\mu C = 70\times 10^{-6}C"

Substituting the numerical values, get:


"\Phi_E = \dfrac{70\times10^{-6}}{8.85\times10^{-12}} \approx 7.91\times10^6V\cdot m"

Answer. 7.91*10^6 V*m.